# Cumulants and ordering of their ratios in 2D Potts models: Lessons for QCD?

### 19 Dec 2023 v1

@@ Abstract
Theoretical considerations suggest an ordering of the ratios of net-baryon number fluctuations in the vicinity of the transition from the low-temperature hadronic phase to the high temperature quark-gluon plasma phase at small values of the baryon chemical potential, $\mu_{B}$, in the QCD phase diagram. The ordering hierarchy is $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, where $\chi_{n}$ is the $n^{\mathrm{th}}$ order cumulant of net-baryon number fluctuation. The STAR experiment observed this hierarchy in the ordering of cumulant ratios of net-proton number (a proxy of net-baryon number) for a range of colliding energies. These inequalities can be tested in spin models by taking the corresponding order parameters in the model as an analog of baryon density. We employed two different models: the two-state and three-state Potts models in two dimensions, which undergo a transition from an ordered phase to a disordered phase at their respective critical temperature. Simulations were performed on square lattices of different sizes using the Wolff algorithm. The cumulants of total magnetization are obtained up to the sixth order in both of these models in a temperature range near their corresponding critical temperatures. With increasing lattice size, height (trough) of the peaks (dips) of the higher-order cumulants appears to increase with the increase in the order of the cumulants. Except in a narrow range above the critical temperature of the three-state Potts model, the complete inequality or its complete reverse is not satisfied in the temperature ranges simulated.
I Introduction
@@
Understanding the phase structure and the nature of phase transitions in the temperature-baryon density (or equivalently the baryon chemical potential, $\mu_{B}$) plane for the strongly interacting matter, known as quantum chromodynamics (QCD) phase diagram, is crucial for unraveling the fundamental properties of matter [1, 2]. While the QCD phase diagram is not well-explored for larger $\mu_{B}$, the region below $\mu_{B}$ = 450 MeV, spanning down to $\mu_{B}=0$, has been extensively studied, both in theory [2, 3] and in experiments [4]. For vanishing and small values of $\mu_{B}$, the transition from the low-temperature hadronic region to the quark-gluon plasma at high temperature is well-known to be a smooth transition (cross over) [5]. At larger values of $\mu_{B}$, however, it is generally expected that a first-order phase transition line exists, which ends in a second-order critical point belonging to the three-dimensional Ising model universality class [6, 7, 8]. Comprehensive studies of net-proton fluctuations have been carried out with the Beam Energy Scan (BES) program of the Relativistic Heavy Ion collider (RHIC) in the STAR experiment to search for the QCD critical point [9, 10]. Recently, an ordering of the cumulant ratios of net-proton number, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, has been reported by the STAR experiment across the energy range 7.7 to 200 GeV [11] where $\chi_{n}$ is the $n^{\mathrm{th}}$ order cumulant of event-by-event net-proton number distribution. This ordering was suggested previously by the Lattice-QCD (LQCD) calculations for the cumulants of net-baryon number fluctuations at small values of $\mu_{\mathrm{B}}$ [12, 13]. Functional renormalization group (FRG) calculations were also found to follow this hierarchy of the inequalities [14]. In contrast, predictions from the Hadron Resonance Gas (HRG) model, utilizing an ideal gas equation of state within a grand canonical ensemble framework, consistently yield positive unity for all cumulant ratios, lacking the observed ordering [15]. One may wonder whether these inequalities are generically valid in the vicinity of any phase transition or even in wider regions surrounding it. In the QCD phase diagram, either near the cross over or near the QCD critical point, if it exists, whether such inequalities exist is an interesting question to ask.
With the advent of lattice gauge theories, the Z(N) spin models have gained significant importance in the study of critical behavior in SU(N) gauge theories at finite temperature. Here, the variable N in general represents the number of degrees of freedom: within SU(N) gauge theories, it denotes the number of colors comprising the gauge group, while in Z(N) spin models, it represents the number of distinct spin states assignable to each lattice site. Among the Z(N) spin models, the three-state Potts model (Z(3)) in three dimensions (3D) holds particular significance as it provides a valuable framework for understanding the phase structures of QCD based on the SU(3) gauge symmetry group. The Polyakov loop, which serves as an order parameter of $\text{SU}(3)$ gauge theories, behaves similarly as spins in the low temperature phase of the 3D three-state Potts model [16, 17]. The line of first order phase transition in QCD phase diagram [18, 19] is closely related to the first order phase transition of 3D three-state Potts model in the absence of an external magnetic field [20, 17]. Addition of a small magnetic field ($h$) leads to a critical point in the Ising universality class for the three-state Potts model in 3D at $h_{c}$ [21], making it a nice toy model to explore the hierarchy of inequalities of the ratios of net-baryon number cumulants. As a first step towards such an investigation, we studied the higher-order cumulants of magnetization in the two-state (Z($2$)) and the three-state (Z($3$)) Potts model on a two-dimensional lattice in this paper. Both are known to have critical points albeit of another universality class. The finite-size scaling of the higher-order cumulants of magnetization were investigated near the critical region in the respective models and the ordering of their ratios were tested.
@@
The paper is structured as follows: in Sec. II, we provide an overview of the $q$-state Potts models, detailing their formalism and the computation of cumulants of magnetization up to $6^{\mathrm{th}}$ order in these models. We also discuss the application of finite size scaling relation, which enables the extraction of exponents associated with the scaling of higher-order cumulants. Section  III presents the methodology employed for simulating the models and outlines the simulation validation process. In Sec. IV, we present the temperature-dependent behavior of cumulants and their ratios. The exponents involved with the finite-size scaling of the higher-order cumulants are estimated in terms of known critical exponents of the model and compared with the numerical results. Finally, Sec. V summarizes the key findings of our study.
II Model and Observables
@@
[…]
$m_{\text{Potts}}=\frac{q}{q-1}\left(\frac{\text{max}(N_{0},N_{1},...,N_{q-1})}{N}-\frac{1}{q}\right)\,,$ (2)
@@
In the following analysis and results, we will be dealing Potts model with both $q$ = 2 and $q$ = 3 states. The two-state Potts model is equivalent to the well-known Ising model for which the exact solution by Onsager gives the critical point at $T_{C}=\frac{2}{\text{ln}(1+\sqrt{2})}\,\mathcal{J}/k_{\text{B}}\approx 2.2692\,\mathcal{J}/k_{\text{B}}$ [23]. The three-state Potts model that belongs to the same universality class as the two-dimensional Ising model, undergoes a second order phase transition at $T_{C}=\frac{1}{\text{ln}(1+\sqrt{3})}\,\mathcal{J}/k_{\text{B}}\approx 0.9950\,\mathcal{J}/k_{\text{B}}$ [22].
II.1 Cumulants
Figure 1: (a) The second order susceptibility of magnetization for the two-dimensional Ising model (equivalent to $q$ = 2 state Potts model) shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). (b) The reduced temperature, $t^{\prime}=T(\chi_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
In this paper, the main objects of interest are the higher-order cumulants (also referred as susceptibilities in the paper), which reflect the higher-order correlations amongst the spins. In spin models, the $p^{\text{th}}$ order susceptibility normalized by the total number of spin sites in the lattice is defined [24] by
$\chi_{p}(T)=\frac{1}{L^{d}\beta^{p}}\left(\frac{\partial^{p}\text{ln}Z}{\partial h^{p}}\right)_{h\to 0}\,,$ (3)
where $\beta=1/(k_{\text{B}}T)$, $T$ is the temperature, $k_{\text{B}}$ is the Boltzmann constant. Here, $Z$ is the partition function of the model given by its corresponding $\mathcal{H}$ as
$Z=\sum_{a}\text{exp}(-\beta\mathcal{H})\,.$ (4)
The sum over $a$ in Eq. 4 means the summation over all possible spin configurations on the lattice. The higher-order spin-spin susceptibilities can be associated with the cumulants of total magnetization ($M$). This is because $Z$ is proportional to the moment generating function [25] of $M$ and therfore one can easily deduce,
@@ II.1 Cumulants
Here, $M$ is referred as magnetization which is equal to the product of the order parameter and the volume of the lattice, i.e., $M=L^{d}m_{\mathrm{Potts}}$. $K_{p}(M)$ is the $p^{\text{th}}$ order cumulant of $M$. The expressions for $\chi_{p}$ (obtained by using the relation of cumulants in terms of central moments [26]) up to $p=6$ is listed below,
@@
where $\mu_{p}=\langle(M-\langle M\rangle)^{p}\rangle$ is the $p^{\text{th}}$ order central moment of the magnetization and the angular brackets, .$\langle$.$\rangle$ . . ⟩ represents average over all possible spin configurations on the lattice. We will be studying these $\chi_{p}$’s as a function of temperature in the critical region.
II.2 Finite size scaling
@@
These exponents for $n$ = 2 to 6 can now be calculated using the respective values of $\gamma$ and $\nu$ for the model under consideration. For the $q=2$ state Potts model on a two-dimensional lattice, we have $\gamma=\frac{7}{4}$ and $\nu=1$ for $q=2$ [29], while $\gamma=\frac{13}{9}$ and $\nu=\frac{5}{6}$ for $q=3$ [22].
Figure 2: (a) The second order susceptibility of magnetization for the two-dimensional $q$ = 3 state Potts model shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). (b) The reduced temperature, $t^{\prime}=T(C_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
III Numerical simulation
@@
Both $q=2$ and $q=3$ state Potts model were simulated on two-dimensional lattices using the Wolff Cluster algorithm [30] at temperatures close to their corresponding critical temperature ($T_{C}$). In this algorithm, clusters of similarly oriented spin sites are determined randomly which are then flipped together to generate new states. This algorithm reduces the effect of critical slowing down near $T_{C}$ with increase in lattice size, i.e. it has a smaller dynamical exponent resulting in more efficient simulations in this region [31]. For $q=2$, square lattices of linear size $L$ = 50, 60, 70, 80 and 90 were used in the simulations and during the simulation $10^{4}$ independent configurations were generated after ensuring thermalization by throwing away sufficient initial configurations. Similarly, for $q=3$, simulations were performed on square lattices with $L$ = 48, 64, 80, 96 and 128 to generate $2\times 10^{4}$ independent configurations after approprite thermalization. In order to obtain independent spin configurations on the lattice, configurations were recorded at the interval of $2\tau$ where $\tau$ is the auto-correlation time that is obtained from the auto-correlation function (of $M$) defined in Eq. 15 using an exponential decay ansatz $\phi(t)\sim\mathrm{exp}\left({-\frac{t}{\tau}}\right)$.
Figure 3: The third and fourth order susceptibilities of magnetization, $\chi_{3}$ and $\chi_{4}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to two-dimensional $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{3}$ and the bottom panel represents $\chi_{4}$ [where $\chi_{3}$ and $\chi_{4}$ divided by a factor $10^{5}$ ($10^{6}$) and $10^{8}$ ($10^{9}$) respectively for $q=2$ ($q=3$)].
IV Results
The top panels of Fig. 1 and Fig. 2 display the second-order susceptibility, denoted as $\chi_{2}$, for the $q=2$ and $q=3$ state Potts model respectively with varying lattice sizes, $L$. In both cases, $\chi_{2}$ exhibits a distinctive peak structure. As may be expected, the locations ($T_{C,L}$) and magnitudes ($\chi_{2}(T_{C,L})$) of these peaks vary with changes in the lattice size. Their scaling behavior can be utilized to extract the critical exponents, specifically $\gamma$ and $\nu$, which will be further elaborated upon in the subsequent section, IV.1.
Figure 4: The fifth and sixth order susceptibilities of magnetization, $\chi_{5}$ and $\chi_{6}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to two-dimensional $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{5}$ and the bottom panel represents $\chi_{6}$ [where $\chi_{5}$ and $\chi_{6}$ divided by a factor $10^{12}$ ($10^{13}$) and $10^{16}$ ($10^{17}$) respectively for $q=2$ ($q=3$)].
IV.1 Critical exponents
The critical exponents, $\gamma$ and $\nu$ can be obtained using the scaling relations, $\chi_{2}(T_{C,L})\propto L^{-\gamma/\nu}$ and $T_{C,L}-T_{C}\propto L^{-1/\nu}$, where $T_{C,L}$ is the psuedo-critical temperature for a given lattice size, and $T_{C}$ represents the critical temperature of the model in the limit $L\to\infty$. The pseudo-critical temperature is an effective critical temperature observed in finite-sized systems, reflecting behavior similar to the critical point in infinite-sized systems. $T_{C,L}$ is defined as the location of the peak in the second order susceptibility, $\chi_{2}$ distribution as shown in Figs. 1 and  2. By fitting $\chi_{2}$ as a Gaussian function of the temperature, $T$, the peak location ($T_{C,L}$, $\chi_{2}(T_{C,L})$) is determined for each lattice size. The obtained values of $T_{C,L}$ and $\chi_{2}(T_{C,L})$ as a function of $L$ are tabulated in Tables 1 and 2 for $q$ = 2 and $q$ = 3 state Potts model respectively. Also, we define the critical region around $T_{C,L}$ for each lattice size by the width of the $\chi_{2}$ distribution peak, which is equal to the $\sigma$ of the fitted Gaussian distribution function, represented here as $\sigma_{C}$. Using the values of $T_{C,L}$, the difference $t^{\prime}=T_{C,L}-T_{C}$ is plotted as a function of $1/L$, and fitted with its scaling relations for both $q=2$ and $q=3$ state Potts model as shown in the bottom panels of Figs. 1 and 2 respectively. Alternatively, $T_{C}$ can be estimated by plotting directly $T_{C,L}$ against $L$ and fitting with $y=A+Bx^{-1/\nu}$. The resulting estimates ($T_{C}=2.2780\pm 0.0089$ for $q=2$ and $T_{C}=0.9953\pm 0.0001$ for $q=3$) are found to be consistent with the known values [23, 22] within uncertainties. For precise measurements of the critical exponents, however we employ the known values of $T_{C}$ in the former method and extracted the coefficient $\nu$. The peak position, $\chi_{2}(T_{C,L})$ as a function of $L$, is also fitted with its scaling relation. Therefore, the extracted values of $\gamma$ and $\nu$ from the fits with the scaling relations are as follows: for $q=2$, $\gamma=1.758\pm 0.060$, $\nu=1.019\pm 0.034$ whereas for $q=3$, $\gamma=1.435\pm 0.014$, $\nu=0.830\pm 0.008$. The excellent agreement observed between the calculated values of $\gamma$ and $\nu$ and their theoretical counterparts affirms the validation of our simulation.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C}$
@@ IV.1 Critical exponents
Table 1: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C}$) of the peak of $\chi_{2}$ distribution as a function of temperature for $q$ = 2 state Potts model. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C}$

### 15 Apr 2026 v2

# Finite-size behavior of higher-order cumulant ratios near criticality in two-dimensional Potts models

@@ Abstract
Theoretical considerations predict a specific hierarchy among ratios of net-baryon number cumulants ($\chi_{n}$, where $n$ is the order of cumulant) in the vicinity of the transition from the low-temperature hadronic phase to the high temperature quark-gluon plasma phase at small baryon chemical potential, $\mu_{\mathrm{B}}$, in the QCD phase diagram. This hierarchy, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, has been observed by the STAR experiment in net-proton number (a proxy of net-baryon number) cumulant ratios over a broad range of collision energies. Motivated by these findings, we investigate whether similar ordering emerges generically in finite statistical systems undergoing second-order phase transitions.
We employ two different spin models: the two-state and three-state Potts models in two dimensions, both exhibiting a transition from an ordered phase to a disordered phase at their respective critical temperatures. Monte Carlo simulations are performed on square lattices of varying sizes using the Wolff cluster algorithm. Cumulants of the total magnetization are calculated up to sixth order in both of these models in a temperature range near their corresponding critical temperatures. Higher-order cumulants exhibit extrema (peaks/troughs) whose magnitudes grow with both cumulant order and lattice size, reflecting enhanced critical fluctuations. Except within a narrow temperature window above the critical temperature, neither the complete hierarchy nor its exact reverse is realized over the studied temperature range in either model.
  • February, 2026
1 Introduction
@@
Understanding the phase structure and the nature of phase transitions in the temperature-baryon density (or equivalently the baryon chemical potential, $\mu_{\mathrm{B}}$) plane for the strongly interacting matter, known as quantum chromodynamics (QCD) phase diagram, is crucial for unraveling the fundamental properties of matter [26, 13]. While this diagram is not well-explored for larger $\mu_{\mathrm{B}}$, the region below $\mu_{\mathrm{B}}$ = 450 MeV ($\mu_{\mathrm{B}}/T_{\mathrm{ch}}\sim$ 2.8–3, where $T_{\mathrm{ch}}$ is the chemical freeze-out temperature inferred from particle yields in heavy-ion collisions using a thermal model), down to $\mu_{\mathrm{B}}=0$, has been extensively studied, both theoretically [13, 9] and experimentally [2]. For vanishing and small $\mu_{\mathrm{B}}$, the transition from the low-temperature hadronic matter to the quark-gluon plasma at high temperature is well-known to be a smooth transition (crossover) [3]. At larger $\mu_{\mathrm{B}}$, however, it is generally expected that a first-order phase transition line exists, which ends in a second-order critical point belonging to the three-dimensional (3d) Ising model universality class [7, 16, 4]. Comprehensive net-proton fluctuation studies have been carried out via the Beam Energy Scan (BES) program of the Relativistic Heavy Ion collider (RHIC) in the STAR experiment to search for the QCD critical point [20, 11].
Recently, an ordering of the cumulant ratios of net-proton number, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, has been reported by the STAR experiment across the energy range 7.7 to 200 GeV [1] (corresponding to an estimated range of $\mu_{\mathrm{B}}\approx$ 20–450 MeV) where $\chi_{n}$ is the $n^{\mathrm{th}}$ order cumulant of event-by-event net-proton number distribution. This ordering was suggested previously by the Lattice-QCD (LQCD) calculations for the cumulants of net-baryon number fluctuations at small values of $\mu_{\mathrm{B}}$ [5, 6]. Functional renormalization group (FRG) calculations were also found to follow this hierarchy of the inequalities [14]. In contrast, predictions from the Hadron Resonance Gas (HRG) model, utilizing an ideal gas equation of state within a grand canonical ensemble framework, consistently yield positive unity for all cumulant ratios, lacking the observed ordering [10]. One may wonder whether these inequalities are generically valid in the vicinity of any phase transition or even in wider regions surrounding it. In the QCD phase diagram, either near the cross over or near the QCD critical point, if it exists, whether such inequalities exist is an interesting question to ask.
Relying on the anticipated universality of the QCD critical point with the 3d Ising model, we investigate in this study whether the observed ordering of higher-order cumulant ratios can arise in simpler two-dimensional (2d) spin models near continuous phase transitions. Quantitative properties of critical behavior, such as critical exponents and scaling functions, depend on dimensionality and are therefore not expected to match those of the QCD critical point. However, qualitative aspects of critical phenomena—such as divergence of the correlation length, enhanced fluctuations of the order parameter, and the scaling structure of higher-order cumulants—are common to systems undergoing second-order phase transitions. In particular, since higher-order susceptibilities correspond to derivatives of the free energy with respect to the external field, their qualitative behavior near criticality can often be explored in simpler statistical systems. The central question we address is therefore whether the ordering of cumulant ratios can emerge as a generic finite-size feature near criticality, or whether it relies on additional ingredients such as dimensionality, explicit symmetry breaking, or the presence of first-order transitions.
The expected QCD phase diagram, with a line of first order phase transitions at large $\mu_{\mathrm{B}}$ which ends in a critical point for smaller $\mu_{\mathrm{B}}$, motivates the use of simplified model system to probe the behavior of these inequalities in finite volumes expected in the heavy-ion collisions. The first order phase transition of 3d three-state Potts model in the absence of an external magnetic field [8, 15] weakens with the addition of a small magnetic field $(h)$ and leads to a critical point in the 3d Ising universality class at $h_{c}$ [17]. Generically, increasing $h$ towards $h_{c}$ should be similar to decreasing $\mu_{\mathrm{B}}$ towards $\mu_{\mathrm{B},c}$ along the first order phase transition line, allowing one to learn more about the inequalities of the cumulant ratios. As a first step towards such an investigation, we have studied the higher-order cumulants of magnetization as a function of temperature for $h=0$ in the two-state ($\mathrm{Z}(2)$) and the three-state ($\mathrm{Z}(3)$) Potts model on 2d lattices in this paper. Both the models are known to have critical points albeit falling under different universality classes. The finite-size scaling of the higher-order cumulants of magnetization were investigated near the critical region in the respective models and the ordering of their ratios were tested. While these studies are ideally needed to be carried out in three dimensions in the presence of nonzero magnetic field, our work nevertheless provides a foundational analysis of the hierarchy of cumulant ratios, which can be generalized to the appropriate model.
It is important to note that the cumulants studied in these spin models differ from those calculated in LQCD. In LQCD at $\mu_{\mathrm{B}}=0$, baryon–antibaryon symmetry implies that all odd-order susceptibilities of the net-baryon number vanish exactly. Consequently, ratios involving odd cumulants become meaningful only when considering finite $\mu_{\mathrm{B}}$, typically through Taylor expansion coefficients of the pressure around $\mu_{\mathrm{B}}=0$ [5]. In contrast, the symmetry structure of the Potts model does not enforce an exact analogue of baryon–antibaryon symmetry for the magnetization variable considered here, and therefore odd-order cumulants of magnetization need not vanish even at $h=0$. Comparison of this study to QCD should therefore be interpreted as qualitative.
@@
The paper is structured as follows: in Sec. 2, we provide an overview of the $q$-state Potts models, detailing their formalism and the computation of cumulants of magnetization up to $6^{\mathrm{th}}$ order in these models. We also discuss the application of finite size scaling relation, which enables the extraction of exponents associated with the scaling of higher-order cumulants. Section  3 presents the methodology employed for simulating the models and outlines the simulation validation process. In Sec. 4, we present the temperature-dependent behavior of cumulants and their ratios. The exponents involved with the finite-size scaling of the higher-order cumulants are estimated in terms of known critical exponents of the model and compared with the numerical results. Finally, Sec. 5 summarizes the key findings of our study.
2 Model and Observables
@@
where $\mathcal{J}$ is a coupling constant in units of energy, determining the interaction strength among neighbouring spins, $h$ is the external magnetic field applied to the model and $\delta$ is the Kronecker delta function. Here, $\sigma_{k}$ can take any integer values in the set {0, 1, 2, …, ($q$ - 1)} and for each $\sigma_{k}$, the spin orientations in the Potts model are represented by $s_{k}=\mathrm{exp}\left(\frac{i2\pi\sigma_{k}}{q}\right)$. The sum is over nearest neighbors $\langle i,j\rangle$. In $d$-dimensions there are 2dd such pairs. We focus on $d=2$ in this work. In the absence of an external magnetic field ($h$ = 0), the model exhibits an exact symmetry represented by the group of permutations $S_{q}$, which encompasses all possible global permutations of the $q$ spin values. This symmetry implies that each spin state is equally likely, and there is no preferred direction for alignment. However, below a critical temperature, the symmetry is spontaneously broken resulting in the alignment of spins along a specific direction and the establishment of an ordered phase. To measure the extent of symmetry breaking in the system, a quantity called the order parameter is used, that provides a measure of the overall alignment of spins in the model. The order parameter for the $q$-state Potts model ($m_{\mathrm{Potts}}$) that we use in this study is given by
$m_{\mathrm{Potts}}=\frac{q}{q-1}\left(\frac{\mathrm{max}(N_{0},N_{1},...,N_{q-1})}{N}-\frac{1}{q}\right)\,$, (2)
@@
In the following analysis and results, we will be dealing Potts model with both $q=2$ and $q=3$ states. The two-state Potts model is equivalent to the well-known Ising model for which the exact solution by Onsager gives the critical point at $T_{C}=\frac{2}{\mathrm{ln}(1+\sqrt{2})}\,\mathcal{J}/k_{\mathrm{B}}\approx 2.2692\,\mathcal{J}/k_{\mathrm{B}}$ [19]. The three-state Potts model that belongs to the same universality class as the 2d Ising model, undergoes a second order phase transition at $T_{C}=\frac{1}{\mathrm{ln}(1+\sqrt{3})}$ $\mathcal{J}/k_{\mathrm{B}}\approx 0.9950\,\mathcal{J}/k_{\mathrm{B}}$ [29].
2.1 Cumulants
In this paper, the main objects of interest are the higher-order cumulants (also referred as susceptibilities in the paper), which reflect the higher-order correlations amongst the spins. In spin models, the $p^{\mathrm{th}}$ order susceptibility normalized by the total number of spin sites in the lattice is defined [21] by
$\chi_{p}(T)=\frac{1}{L^{d}\beta^{p}}\left(\frac{\partial^{p}\mathrm{ln}Z}{\partial h^{p}}\right)_{h\to 0}\,$, (3)
where $\beta=1/(k_{\mathrm{B}}T)$, $T$ is the temperature, $k_{\mathrm{B}}$ is the Boltzmann constant. Here, $Z$ is the partition function of the model given by its corresponding $\mathcal{H}$ as
$Z=\sum_{a}\mathrm{exp}(-\beta\mathcal{H})\,.$ (4)
The sum over $a$ in Eq. 4 means the summation over all possible spin configurations on the lattice. The higher-order spin-spin susceptibilities can be associated with the cumulants of total magnetization ($M$). This is because $Z$ is proportional to the moment generating function [25] of $M$ and therefore one can easily deduce,
@@ 2.1 Cumulants
Here, $M$ is referred as magnetization which is equal to the product of the order parameter and the volume of the lattice, i.e., $M=L^{d}m_{\mathrm{Potts}}$. $K_{p}(M)$ is the $p^{\mathrm{th}}$ order cumulant of $M$. The expressions for $\chi_{p}$ (obtained by using the relation of cumulants in terms of central moments [27]) up to $p=6$ is listed below,
@@
where $\mu_{p}=\langle(M-\langle M\rangle)^{p}\rangle$ is the $p^{\mathrm{th}}$ order central moment of the magnetization and the angular brackets, $\langle..\rangle$ represents average over all possible spin configurations on the lattice. We will be studying these $\chi_{p}$’s as a function of temperature in the critical region.
2.2 Finite size scaling
@@
Using Eq. (13), ratios of two susceptibilities can be expressed as
$\frac{\chi_{m}}{\chi_{n}}=C_{2}^{m-n}L^{\frac{(m-n)\Delta}{\nu}}\frac{\partial^{m}Y/\partial h_{r}^{m}}{\partial^{n}Y/\partial h_{r}^{n}}\bigg|_{h_{r}=0}.$ (14)
While the leading lattice-size dependence is reduced in such ratios compared to individual susceptibilities, they retain contributions from the non-universal amplitude $C_{2}$ as well as from derivatives of the scaling function $Y$. Since scaling theory does not constrain the relative magnitudes or signs of the derivatives of $Y$, it does not imply any particular ordering among the cumulant ratios. Consequently, scaling theory by itself does not enforce any hierarchy among the cumulant ratios; their ordering depends on the detailed form of the scaling functions.
@@
These exponents for $n$ = 2 to 6 can now be calculated using the respective values of $\gamma$ and $\nu$ for the model under consideration. For the $q=2$ state Potts model on a 2d lattice, we have $\gamma=\frac{7}{4}$ and $\nu=1$ for $q=2$ [23], while $\gamma=\frac{13}{9}$ and $\nu=\frac{5}{6}$ for $q=3$ [29].
3 Numerical simulation
@@
Both $q=2$ and $q=3$ state Potts models were simulated on 2d lattices using the Wolff Cluster algorithm in a region around the corresponding critical temperature ($T_{C}$) [28]. All simulations were performed in the absence of an external magnetic field, i.e., hh(or $h_{r}$) = 0. In the Wolff algorithm, clusters of similarly oriented spin sites are determined randomly which are then flipped together to generate new states. This algorithm reduces the effect of critical slowing down near $T_{C}$ with increase in lattice size, i.e. it has a smaller dynamical exponent resulting in more efficient simulations in this region [18]. We have used periodic boundary conditions for the finite lattices, which are also commonly employed in simulations of LQCD at finite temperature and density [12]. For $q=2$, square lattices of linear size $L$ == 50, 60, 70, 80 and 90 were used in the simulations and during the simulation $10^{4}$ independent configurations were generated after ensuring thermalization by throwing away sufficient initial configurations. Similarly, for $q=3$, simulations were performed on square lattices with $L$ == 48, 64, 80, 96 and 128 to generate $2\times 10^{4}$ independent configurations after appropriate thermalization. Simulations were conducted for the temperature range $0.95T_{C}\leq T\leq 1.05T_{C}$, and the corresponding results are presented. Tables 1 and 2 show that this range adequately covers the estimated critical region for all lattice sizes examined.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C,L}$
@@
Table 1: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C,L}$) of the peak of $\chi_{2}$ distribution on an $L^{2}$ lattice as a function of temperature for $q=2$ state Potts model. The quoted values after $\pm$’ represent statistical uncertainties derived from the fit.
Lattice size $T_{C,L}$ $\chi_{2}(T_{C,L})$ $\sigma_{C,L}$